1) What are the three different kinds of rate equations?

2) How do these rate equations differ from one another?

3) How do you solve for these rate equation?


1
Expert's answer
2021-05-28T00:56:22-0400

1) There are thee types of the rate equations:


i. Zero order reaction : For zero-order reactions, the reaction rate is independent of the concentration of a reactant, so that changing its concentration has no effect on the speed of the reaction. Thus, the concentration changes linearly with time.


ii. First order reaction : first order reaction depends on the concentration of only one reactant (a uni molecular reaction). Other reactants can be present, but each will be zero order.


iii. Second order reaction : A reaction is said to be second order when the overall order is two.


2.Difference in the three types of rate equations:


In zero order reaction the the Rate of the equation is independent to the concentration of the reactant.

In First order reaction the Rate of the equation depends on the single power of the reactant.

In Second order reaction the Rate of the equation is proportional to the square of the reactant.


3.Solution of rate equations :


i. Zero order reaction : "[C]=[C_0] -kt"


ii. First order reaction : "[C]=[C_0]\times e^{-kt}"


iii. Second order reaction : "\dfrac{1}{[C]} = \dfrac{1}{[C_0]} -kt"


where: "[C]" = concentration of reactant at time 't'

"[C_0]" = Initial concentration of reactant.

"k" = Rate constant of the reaction.